OpenStudy (anonymous):

Which is an example of the commutative property? A. –13 + 27 = 27 + (–13) B. –21 + (11 + 13) = (–21 + 11) + 13 C. 4(–3 + 5) = 4(–3) + 4 ∙ 5 D. –(–8 + 3) = 8 + (–3)

OpenStudy (mathstudent55):

Commutative: a + b = b + a (change of order) Associative: a + (b + c) = (a + b) + c (change of grouping)

OpenStudy (anonymous):

A does commutative property allows you to swap numbers and have it sstill mean the same

OpenStudy (anonymous):

Thanks! :) Both very helpful answers!